A 10m long clothesline


1
Expert's answer
2021-06-16T15:38:22-0400
"\tan \theta=\frac{d_e}{d}\\\tan 1=\frac{W}{5}\\W=0.0873\ m"

Hence, the cloth line will be x = 0.0873 m from the original length of the clothesline.